Grade 12th passGeneral Physics

A 24kg block resting on a floor has a rope tied to its top . the maximum tension that rope can with stand without breaking is 310N .the minimum time in the block can be lifted a vertical distance of 4.6m by pulling on the rope

Ratchana Kumar
9 Years agoGrade 12th pass

1 Answer

Vikas TU
8 Years ago
Dear Student,
For minimum time the acceration should be max
the condition thus becomes 24g+24a=310
(24a is the pseudo force that will act downwards.)
=>a=35/12 m/s
Cheers!!
Regards,
Vikas (B. Tech. 4th year
Thapar University)